OpenStudy (anonymous):

how do you convert 11/32 to a percent

OpenStudy (mertsj):

Divide 11.00000 by 32 and then move the decimal point 2 places to the right.

OpenStudy (anonymous):

11/32=x/100 32x=11(100) x=11(100)/32 - put that on your calculator